Colors Fonts Set as Default Document Formatting Answer the following questions: Question 1. [3+5=8 Marks] Random samples have I need the answer of part B   
Answer #1

solution Sumple from population - 1 (X) 14 Sample from 2 (X2) population 12 15 lo 15 14 Jo 12 10 13 The above data by calculaet t(0.0512) = 1.782 (31-1) 5,² + (02-1) 5₂² nith2-2 (6) (1.35)² + (6) (0-62) 2 Sp = 1.4911 CI = (2.57 + (782)(19911) V7th =(D 2010 The null and alternative hypothesis is the Horrel, -4250 VS Hill-e1280 Test statistics X-82 To= Sp n + 1 (14.14-11:57)Critical value tegnitiz-2) = 4(0777-2) =t(0.05,12) =1782 Conclusion since )To) - 3-2245 >t(0.05, 12) = 1.782 . We reject to